OpenStudy (anonymous):

write the mixed number one and three fourths as an improper fraction then multiply the improper fraction by one half what is the product

OpenStudy (anonymous):

7/4,7/8

OpenStudy (anonymous):

1 + (3/4) = (4/4)+(3/4) = (7/4) the improper fraction as the top is largest. remember product is multiply so need 7/4 multiply by 1/2 (7/4) (1/2) = (7 x 1) /(4 x 2) = (7/8)